if a triangle has side lengths 5 cm, 7 cm, and 9cm does it form a right triangle ? explain.
Found 2 solutions by stanbon, PkGeci:
Answer by stanbon(75887)

if a triangle has side lengths 5 cm, 7 cm, and 9cm does it form a right triangle ? explain.
--------
right triangle if 5^2 + 7^2 = 9^2
Checking:
25 + 49 = 81
---
Answer No, so not a right triangle.
Cheers,
Stan H.
===============

Answer by PkGeci(6)

Pythagorean theorem states that a^2+b^2 = c^2
therefore the triangle with sides 5 cm, 7 cm and 9 cm do not form a right triangle because 5^2 + 7^2 does not equal 9^2